Question 1

(a)The law of demand is an empty box devoid of real world applications, discuss.


(b) Are there instances when the law of demand and supply suspended, in those situations, the consumer is likely to buy more as price rises and the producer sells more when price falls, discuss.


(c) Given the following demand model:

Qd = 50 – 5P; Qs = -30 + 10P;

Qd = Qs.

Solve for the equilibrium price and quantity demanded.




1
Expert's answer
2020-11-19T06:38:18-0500

a)This is because sometimes rising incomes tend to increase demand for normal economic goods, as people are willing to spend more. Additionally, the availability of close substitute products that compete with a given economic good will tend to reduce demand for that good, since they can satisfy the same kinds of consumer wants and needs


b) Prestigious goods, Veblen goods, Giffen goods, and stock exchanges are the major good s that violates the law of demand. For instance, even if the price for prestigious goods goes up, the demand will still be the same.



c)"Q_d=Q_s"


Therefore, "50 \u2013 5P = -30 + 10P;"


"-15p=-80"


"p=5.33"


substituting p in Qd = 50 – 5P:


"Qd=50-5(5.33)=23.35"


Equilibrium quantity is "23.35"